OpenStudy (anonymous):

I have a PS (problem solving) and i need to figure out an equation to get my answer, but I'm stuck. it sounds somewhat like this: 9 over S = 12 over (s-2) minus 3 over (s-6) if anyone can write this out and solve it i would be very happy. Thank you

OpenStudy (anonymous):

This is a 7th grade problem!! Please help me!! its due tomorrow :(

OpenStudy (dumbcow):

\[\frac{9}{s} = \frac{12}{s-2} -\frac{3}{s-6}\] combine fractions on right side \[\frac{9}{s} = \frac{12(s-6)-3(s-2)}{(s-2)(s-6)}\] distribute and add like terms on top then cross multiply \[9(s-2)(s-6) = s(9s-66)\] expand and set =0 s=18

OpenStudy (anonymous):

Thank you so much! :)

OpenStudy (dumbcow):

your welcome